OpenStudy (anonymous):

help !Solve: 4(x − 1) = 2(x + 2) + 6 14 8 7 28

OpenStudy (anonymous):

we have tried to help you. Me and @precal. You want answers. we are here to help.

OpenStudy (anonymous):

i got 14

OpenStudy (anonymous):

is that right??

OpenStudy (precal):

show us your work, use draw.........

OpenStudy (anonymous):

7

OpenStudy (precal):

we can help you if you are incorrect and help guide you learn the mathematics along the way

OpenStudy (anonymous):

@ASAAD123 dont just give answers

OpenStudy (anonymous):

4x-4=2x+4+6 4x-2x=4+6+4 2x=14 x=7

OpenStudy (anonymous):

but there is x on both sides'

OpenStudy (precal):

you can move the x to whichever side you want ASAAD123 decided to move it to the left side

OpenStudy (precal):